In a monopoly market, if the firm's market demand function is reflected at P = 11,100 - 30 Q, while the company's total cost function is TC = 400.000 + 300Q – 30 Q2+ Q3

Determine:

a. the output level and selling price per unit that maximizes the company’s profit?

b. maximum total profit?

c. quantity at the socially optimum price (P = MC)?

d. quantity at the fair return price (P = ATC)?
Answer #1

Solution Demand Function is. P. 11,100 - 300 a ) 70 = 400.000 300-300?Q? TR =PQ FR = L1H100 - 300) Q TR 11.200Q. -30Q2 s T .T(b) Substitude into the total profit cocation: H, 1000 -30Q² =400,000 - 3000+ 30002-03 TT - 11,100 (3,60o) -363.600)2 - 400.0- 30+ 1900 +19.9600 -6 = 30 + 366-249 -6 Q 265.208. (d) Quality at the fair retum Price P = ATC livom inftrmation is: Pa tt,